OpenStudy (anonymous):

What is the simplified form of the expression square root of 1 over 169? (1 point) 1 over 338 one thirteenth 13 2 over 169

OpenStudy (anonymous):

\[\sqrt{\frac{ 1 }{ 169 }} = \frac{ \sqrt{1} }{ \sqrt{169} }\]

OpenStudy (anonymous):

the square root of 1 is 1 so \[\frac{ 1 }{\sqrt{169} }\]

OpenStudy (anonymous):

and the square root of 169 is 13, that's just something you end up remembering

OpenStudy (anonymous):

so the answer is \[\frac{ 1 }{ 13 }\]

OpenStudy (anonymous):

the trick to these is knowing that the square root of a fraction: \[\sqrt{\frac{ a }{ b }}\]

OpenStudy (anonymous):

can be written as a fraction of square roots: \[\frac{ \sqrt{a} }{ \sqrt{b} }\]